OpenStudy (anonymous):

f(x) = the quantity x minus seven divided by the quantity x plus three. and g(x) = quantity negative three x minus seven divided by quantity x minus one.

OpenStudy (anonymous):

\[f(x)=\frac{x-7}{x+3}\]\[g(x)=\frac{-3x-7}{x-1}\]